# Continuous-Time State Estimation Methods in Robotics: A Survey

### v1

@@ IV-B Temporal Splines
IV-B1 Origins & Early Mathematical Formulations
Splines have received significant attention in robotics as a CT representation of the state. They were first formulated as piecewise-polynomial by Schoenberg in 1946 [178], inspired by ‘lofting’ where thin wooden strips (called splines) were used for aircraft and ship design, including for construction from templates during World War II [179]. While various spline types have been explored in different robotics applications, B-splines, short for basis splines, are most popular due to their smoothness properties. The mathematics of B-splines was pioneered by Schoenberg, de Boor, and others, with the de Boor-Cox recurrence formula [49, 50] a famous early formulation. From this, Qin [52] was the first to derive a general matrix representation for B-splines. Later, splines were presented as a special case of Temporal Basis Functions (TBFs) where the bases are polynomials [20, 134]. As an alternative to polynomials, wavelet basis functions [48, 127] have been proposed in a hierarchical scheme to account for the varying richness of motion. Sinusoidal temporal bases have also been used in conjunction with polynomials for the problem of certifiable range-based localization [158] or object tracking [149, 150, 151].
@@ IV-B3 Choice of Pose Representation
For 3D pose estimation using B-splines, a choice exists between the joint () and split ( or ) representations. Ovrén and Forssén [157] scrutinize this choice for robotic applications and advocate for the split representation in most cases. They showed that under the representation, translation, and orientation are coupled in screw motion, where implicitly “the pose is manipulated by an internal force and torque”, an action called a wrench [180], which may be appropriate in some contexts such as for the trajectories of cameras mounted on a robot arm. In contrast, they argue that the split representation implicitly assumes that an external force and torque actuate the pose. Furthermore, they demonstrate that the analytic acceleration of the joint spline is not the kinematic body acceleration in general. Besides the theoretical justifications, they experimentally show that the split representation converges faster and more accurately for their RS Visual-Inertial SLAM system. More recently, Tirado and Civera [55] make a similar argument; however, they advocate for the joint representation as the general case for rigid-body motion. Hug and Chli [31] prefer the split representation, emphasizing the increased computational effort required in the representation. For orientation, Haarbach et al. [46] and Li et al. [56] advocate for quaternions () over rotation matrices (). This can be attributed to their space efficiency, requiring four variables instead of nine, and their relative computational efficiency. For the same reasons, Haarbach et al. [46] advocate for the unit dual quaternion over for B-spline pose interpolation.
IV-B4 Modern Mathematical Advancements
@@
Sommer et al. [53] devise a computationally efficient method for computing Lie group B-spline derivatives, requiring instead of matrix operations for a spline of order . They further derive a recursive strategy to compute analytic Jacobians for the B-spline and its first and second derivatives in time. This was extended to other Lie groups [55, 56] with Tirado and Civera [55] demonstrating significant computational savings over automatic or numerical differentiation. Finally, Johnson et al. [22, Appendix E] provide analytic Jacobians of general Lie group splines up to the second derivative. Hug and Chli [31] contribute analytic Jacobians for the knots, suggesting the possibility that they could be jointly optimized in future work. Recently, Hug et al. [176] use SymForce [181] to obtain analytic B-spline implementations that are much faster than those derived by Sommer et al. [53]. Persson et al. [45] provide a number of novel theoretical contributions: i) constant-velocity extrapolation through new basis functions, ii) a necessary condition on knot spacing for quaternion splines to express a maximum angular velocity, iii) a simple method to derive and compute integrals of spline derivative functions, iv) a theoretical and experimental investigation of different spline regularizers, and v) an approach for how these integral-type costs should be sampled. Spline regularization has been effective in reducing oscillations [20, 18] and enforcing non-holonomic constraints [163]. Additionally, Cioffi et al. [19] demonstrated with batch optimization that B-splines are more accurate than DT for trajectory estimation, especially when estimating for time offsets between sensors.
@@ IV-B7 Localization, Odometry & SLAM
Sheehan et al. [139] use a Catmull-Rom spline for LiDAR localization, and the recent work by Li et al. [56] uses B-splines for UWB-inertial localization. B-splines have been used to process high-rate event camera measurements, which are particularly hard to aggregate for DT systems. This includes event [144] and event-inertial localization [155], and event VO [167]. Leveraging prior work [74], Nguyen et al. [174] propose an iterative, parallelized, on-manifold linear solver for real-time surfel-based LIO. Lv et al. [175] use a ‘spot uncertainty model’ to weight point-to-plane residuals in their LO system. Ng et al. [161] apply B-splines to RADAR-inertial odometry, using radial velocity measurements from multiple radars to form ego-velocity residuals. Adopting uniform cubic B-splines for the trajectory, Li et al. [177] jointly optimize LiDAR point-to-plane, RADAR Doppler velocity, and inertial residuals for multi-sensor odometry. B-splines have also been used as the CT formulation for robot trajectories in many SLAM works. This includes RS visual [135, 142, 147], RS visual-inertial [168], visual-inertial [157, 31, 164], multi-camera [162], LiDAR [154], and LiDAR-visual-inertial [170, 171] SLAM.
@@ IV-B8 Sensor Calibration
TBFs and B-splines have been used for sensor calibration, including estimation of the line delay of a RS camera [138], visual-inertial [18] and LiDAR-visual-inertial [140] extrinsic calibration, and comprehensive simultaneous intrinsic and extrinsic calibration for LiDAR-visual-inertial [145], multi-IMU visual-inertial [146], and visual-inertial [152, 45] configurations. More recently, uniform quintic B-splines have been applied to RS visual-inertial calibration [166], and uniform cubic B-splines to targetless LiDAR-inertial calibration in structured environments [159, 169], both covering many intrinsic and extrinsic parameters.
@@ IV-B9 Online Optimization
While many perform batch optimization, some works operate online using sliding-window optimization (a.k.a. fixed-lag smoothing). The first to do so only used linear splines (equivalent to LI) [99, 100, 101, 104], but Zlot and Bosse [100] also did with cubic B-splines. Persson et al. [45] were among the first to use windowed optimization for real-time odometry. Lang et al. [168] performed RS visual-inertial SLAM in a sliding-window optimization scheme with B-splines but could not run in real-time. Lv et al. [170] extended this method to a real-time LiDAR-Visual-Inertial system, developing marginalization strategies for B-spline odometry and SLAM problems. More recent spline-based odometry systems have continued to develop marginalization strategies [171, 175, 177], with Lang et al. [171] to first to do so with non-uniform splines. Importantly, the use of per-point residuals in these works allows MDC to occur simultaneously to state estimation [170, 171, 175, 177]. Quenzel and Behnke [54], Mo and Sattar [165] who incorporate inertial measurements through cubic splines into DSO [182], and Tirado and Civera [55] who employ non-uniform cubic B-splines for object tracking, also employ sliding-window optimization. In contrast to these optimization approaches, Li [172] devises a probabilistic recursive filter that maintains the control points for a single cubic B-spline segment in its state and can thus run in real-time with uncertainty estimation. Hug et al. [176] take a different approach, proposing a Gaussian Belief Propagation (GBP) optimization framework called Hyperion, where factor graph optimization is achieved through efficient message passing between nodes and factors. While usually slower than existing Nonlinear Least Squares (NLLS) solvers such as Ceres [183], GBP computes variable (inverse) uncertainty during optimization instead of as a post-optimization procedure, allowing strategies for targeted updating (e.g., of non-converged nodes), and is inherently scalable to multi-agent systems.
@@ IV-C1 Early Works
The use of GPs for CT state estimation was first realized by Tong et al. [57, 47, 217], where they presented Gaussian Process Gauss-Newton (GPGN), a nonlinear batch optimization method, and applied it to 2D range-based localization [57] and SLAM [47]. In these early formulations, the connection with TBFs was made explicit in the derivation, with the transition from parametric to non-parametric achieved through the kernel trick. Tong et al. [184] subsequently extended the GPGN method to 3D for their laser-based visual SLAM system. GPGN preceded the works of Barfoot et al. [34] and Anderson et al. [35], who demonstrated that GPs using the LTV-SDE class of prior have exactly sparse inverse kernel matrices, enabling efficient regression and interpolation. The computation of the inverse kernel matrix is reduced from cubic to linear time, and interpolation can be computed in constant time. When using this prior with a linear measurement model, this approach is shown to be exactly equivalent to discrete-time smoothing when evaluated at the measurement times with identical computational complexity [35, 24]. Solving the SLAM problem using this continuous-time method was called Simultaneous Trajectory Estimation And Mapping (STEAM), with Anderson and Barfoot [59] extending the approach to poses using local LTV-SDEs (Section III-C1). Nonlinear models are supported in these works through linearization and iterative optimization.
@@ IV-C2 Optimization
Exploiting the sparsity of the formulation, the STEAM problem can be modeled as a factor graph, with binary GP prior factors connecting adjacent states [35, 59]. Yan et al. [185, 186] used this to develop a GP version of the incremental optimizer iSAM2 [218], and with iGPMP2 [58] applied this incremental GP approach to motion-planning problems. Sparse interpolated measurement factors allow measurements to be added between the explicitly estimated states [47, 35, 186, 36], and can be readily used within an incremental Bayes tree optimization scheme [186]. Dong et al. [60, 61] provided a formulation for general matrix Lie groups before integrating it into their Simultaneous Trajectory Estimation And Planning (STEAP) framework [188, 189].
Exploring a different direction, Barfoot et al. [196] propose Exactly Sparse Gaussian Variational Inference (ESGVI) to fit a Gaussian distribution to the full Bayesian posterior of the problem states. This approximation of the posterior ought to be more representative than the Laplace approximation obtained from MAP estimation (Equation 6). They devise a Newton-style iterative solver that avoids the need for analytic derivatives through cubature rules and exploit the inherent sparsity of the problem structure to achieve the same computational complexity as MAP estimation. Wong et al. [197] then extend this approach for hyperparameter learning with an Expectation Maximization (EM) optimization framework.
@@ IV-C3 Applications using the WNOA Prior
The GN TGP approach [35, 59, 60] has been applied in the context of teach and repeat, such as in Visual Odometry (VO) for UAVs [192, 193], or LiDAR and RADAR odometry [206, 213]. Liu et al. [207] and Wang and Gammell [208] have proposed incremental event-based VO pipelines based on this CT representation, both employing the GN algorithm in the underlying optimization. Li et al. [202] demonstrated LiDAR-inertial extrinsic calibration and localization in structured environments. Incremental optimizers based on the Bayes tree [218] have been used in SLAM for agricultural monitoring [187] and inertial-GNSS odometry [204]. Two recent works [205, 209] have applied TGPs to odometry with Frequency-Modulated Continuous-Wave (FMCW) LiDARs. Rather than aggregating LiDAR points for registration or feature extraction, Wu et al. [205] formulate point-to-plane and radial-velocity residuals for each point. By contrast, Yoon et al. [209] combine the radial-velocity residuals with a gyroscope for linear estimation of the total DoF ego-velocity, which they integrate into poses. Judd and Gammell [195, 194] address the multimotion estimation problem with TGPs in their occlusion-aware VO pipeline, estimating for the ego-motion and trajectories of points on dynamic objects.
@@ IV-C4 Applications using Other Priors
The aforementioned works all use the WNOA prior. Tang et al. [62] propose the ‘constant-acceleration’ White-Noise-On-Jerk (WNOJ) model instead. They achieve higher accuracies in some cases and prove that when the chosen prior does not match the motion, the state estimate will be biased along certain DoF. The WNOJ prior has since been used for online GNSS multi-sensor odometry [203, 216]. Recently, Zheng and Zhu [212] propose multi-sensor LIO using WNOJ for position, WNOA for orientation, and White-Noise-On-Velocity (WNOV) (a.k.a. ‘random-walk’) for IMU biases, and achieve state-of-the-art results. As in Wu et al. [205] and their previous work [30], they avoid measurement aggregation (e.g., LiDAR registration, IMU pre-integration) while retaining real-time performance through the split pose representation and efficient map management. Burnett et al. [214] also utilize the WNOV prior for IMU biases, demonstrating effective LiDAR-inertial and RADAR-inertial odometry on various datasets. Judd and Gammell [200, 201] compare the WNOJ and WNOA priors in their Multimotion Visual Odometry (MVO) pipeline [194, 195]. Johnson et al. [22] recently were the first to compare TGPs and B-splines in visual-inertial localization experiments. They compared WNOJ TGPs to cubic and quintic B-splines, both with and without motion prior factors, and for both the split and joint pose representations. They conclude that when the same motion models and measurements are used and the degree-of-differentiability matches, temporal splines and GPs are similar in accuracy and solve time. Furthermore, they show that motion prior factors act as regularizers, preventing overfitting to noisy measurements. However, the benefit of these is significantly diminished when IMU is available. Zheng and Zhu [30] and Johnson et al. [22] investigate the efficacy of the split pose representation and conclude its primary benefit is the reduction of computation time. Most recently, Nguyen et al. [63] open-sourced their split pose WNOJ implementation and provide localization and calibration examples for a variety of multi-sensor combinations. They further provide a generalization of White-Noise-On-Derivative (WNOD) priors for any order and analytic Jacobians for all their residuals. Finally, Wong et al. [198] investigate the Singer prior, which represents latent accelerations as a GP with a Matérn kernel. They argue that the data-driven optimization of an additional parameter enables a better representation of the robot’s motion.

### v2

@@ IV-B Temporal Splines
IV-B1 Origins & Early Formulations
Splines have received significant attention in robotics as a CT representation of the state. They were first formulated as piecewise-polynomial by Schoenberg in 1946 [188], inspired by ‘lofting’ where thin wooden strips (called splines) were used for aircraft and ship design, including for construction from templates during World War II [189]. While various spline types have been explored in different robotics applications, B-splines, short for basis splines, are most popular due to their smoothness properties. Their mathematical foundations were pioneered by Schoenberg, de Boor, and others, with the de Boor-Cox recurrence formula [51, 52] a famous early formulation. From this, Qin [54] was the first to derive a general matrix representation for B-splines. Later, splines were presented as a special case of TBFs where the bases are polynomials [20, 141]. As an alternative to polynomials, wavelet basis functions [50, 134] have been proposed in a hierarchical scheme to account for the varying richness of motion. Sinusoidal temporal bases have also been used in conjunction with polynomials for certifiable range-based localization [165] and object tracking [156, 157, 158].
@@ IV-B3 Choice of Pose Representation
For 3D pose estimation using B-splines, a choice exists between the joint () and split ( or ) representations. Ovrén and Forssén [164] scrutinize this choice for robotic applications and advocate for the split representation in most cases. They showed that under the representation, translation, and orientation are coupled in screw motion, where implicitly “the pose is manipulated by an internal force and torque”, an action called a wrench [190], which may be appropriate in some contexts such as robot arm end effector trajectories. In contrast, they argue that the split representation implicitly assumes that an external force and torque actuate the pose. Furthermore, they demonstrate that the analytic acceleration of the joint spline is not the kinematic body acceleration in general. Besides the theoretical justifications, they experimentally show that the split representation converges faster and more accurately for their RS Visual-Inertial SLAM system. Tirado and Civera [57] make a similar argument; however, they advocate for the joint representation as the general case for rigid-body motion. Hug and Chli [31] prefer the split representation, emphasizing the increased computational effort required in the representation. For orientation, Haarbach et al. [48] and Li et al. [58] advocate for quaternions () over rotation matrices (). This can be attributed to their space efficiency, requiring four variables instead of nine, and their relative computational efficiency. For the same reasons, Haarbach et al. [48] advocate for the unit dual quaternion over for B-spline pose interpolation.
IV-B4 Modern Advancements
@@
Sommer et al. [55] devise a efficient method for computing Lie group B-spline derivatives, requiring instead of matrix operations for a spline of order . They further derive a recursive strategy to compute analytic Jacobians for the B-spline and its first and second derivatives in time. This is extended to other Lie groups [57, 58] with significant computational savings demonstrated over automatic or numerical differentiation [57]. Finally, Johnson et al. [22, Appendix E] provide analytic Jacobians of general Lie group splines up to the second derivative. Hug and Chli [31] contribute analytic Jacobians for the knots, suggesting the possibility that they could be jointly optimized in future work. Recently, Hug et al. [184] use SymForce [191] to obtain analytic B-spline implementations that are much faster than those derived by Sommer et al. [55]. Persson et al. [47] provide a number of novel theoretical contributions: i) constant-velocity extrapolation through new basis functions, ii) a necessary condition on knot spacing for quaternion splines to express a maximum angular velocity, iii) a simple method to derive and compute integrals of spline derivative functions, iv) a theoretical and experimental investigation of different spline regularizers, and v) an approach for how these integral-type costs should be sampled. Spline regularization has been effective in reducing oscillations [20, 18] and enforcing non-holonomic constraints [170]. Additionally, Cioffi et al. [19] demonstrated with batch optimization that B-splines are more accurate than DT for trajectory estimation, especially when estimating for time offsets between sensors.
@@ IV-B7 Localization, Odometry & SLAM
Sheehan et al. [146] use a Catmull-Rom spline for LiDAR localization, and Li et al. [58] use B-splines for UWB-inertial localization. B-splines have been used to process high-rate event camera measurements, which are particularly hard to aggregate for DT systems. This includes event [151] and event-inertial localization [162], event VO [174], and map-free event-inertial ego-velocity estimation [181]. Leveraging prior work [77], Nguyen et al. [182] propose an iterative, parallelized, on-manifold linear solver for real-time surfel-based LIO. Lv et al. [183] use a ‘spot uncertainty model’ to weight point-to-plane residuals in their LO system. Ng et al. [168] apply B-splines to RADAR-inertial odometry, using radial velocity measurements from multiple radars to form ego-velocity residuals. Adopting uniform cubic B-splines for the trajectory, Li et al. [185] jointly optimize LiDAR point-to-plane, RADAR Doppler velocity, and inertial residuals for multi-sensor odometry. B-splines have also been used as the CT trajectory formulation in RS visual [142, 149, 154], RS visual-inertial [175], visual-inertial [164, 31, 171], multi-camera [169], LiDAR [161], and LiDAR-visual-inertial [177, 178, 179] SLAM.
@@ IV-B8 Sensor Calibration
TBFs and B-splines have been used for sensor calibration, including estimation of RS camera line delay [145, 173, 175], visual-inertial [18] and LiDAR-visual-inertial [147, 177] extrinsic calibration, and comprehensive simultaneous intrinsic and extrinsic calibration for LiDAR-visual-inertial [152], multi-IMU visual-inertial [153], and visual-inertial [159, 47] configurations. Uniform quintic B-splines have been applied to RS visual-inertial calibration [173], and uniform cubic B-splines to targetless LiDAR-inertial calibration in structured environments [166, 176].
@@ IV-B9 Online Optimization
While many perform batch optimization, some works operate online using sliding-window optimization (a.k.a. fixed-lag smoothing). Early works use linear splines (equivalent to LI) [102, 103, 104, 107], but Zlot and Bosse [103] also did with cubic B-splines. Persson et al. [47] were among the first to use windowed optimization for real-time odometry. Lang et al. [175] develop B-spline marginalization strategies for RS visual-inertial fixed-lag smoothing, refined by Lv et al. [177] to achieve real-time sliding-window LiDAR-visual-inertial SLAM. More recent spline-based odometry systems have continued to develop marginalization strategies [178, 183, 185], with Lang et al. [178] to first to do so for non-uniform splines. Importantly, the use of per-point residuals in these works allows MDC to occur simultaneously to state estimation [177, 178, 183, 185]. Quenzel and Behnke [56], Mo and Sattar [172] who incorporate inertial measurements through cubic splines, and Tirado and Civera [57] who employ non-uniform cubic B-splines for object tracking, also employ sliding-window optimization. In contrast to these optimization approaches, Li [45] devises a probabilistic recursive filter that maintains the control points for a single cubic B-spline segment in its state and can thus run in real-time with uncertainty estimation. Hug et al. [184] take a different approach, proposing a Gaussian Belief Propagation (GBP) optimization framework called Hyperion, where factor graph optimization is achieved through efficient message passing between nodes and factors. While usually slower than existing Nonlinear Least Squares (NLLS) solvers such as Ceres [192], GBP computes variable (inverse) uncertainty during optimization instead of as a post-optimization procedure, allowing strategies for targeted updating (e.g., of non-converged nodes), and is inherently scalable to multi-agent systems.
@@
| Peršić et al. [208] | 2021 | Spatiotemporal Calibration from Object Tracking | - | WNOJ TGPs | ||||||||||||||||
@@
| Burnett et al. [228] | 2024 | LIO | - | Singer & WNOV TGPs | ||||||||||||||||
| Shen et al. [46] | 2024 | Multi-LiDAR Odometry | - | Constant Angular Velocity & Acceleration TGP | ||||||||||||||||
@@ IV-C1 Early Works
The use of GPs for CT state estimation was first realized by Tong et al. [59, 49, 229], where they presented Gaussian Process Gauss-Newton (GPGN), a nonlinear batch optimization method, and applied it to 2D range-based localization [59] and SLAM [49]. In these early formulations, the connection with TBFs was made explicit in the derivation, with the transition from parametric to non-parametric achieved through the kernel trick. Tong et al. [193] subsequently extended GPGN to 3D for their laser-based visual SLAM system. Crucially, Barfoot et al. [34] and Anderson et al. [35] demonstrated that GPs using the LTV-SDE class of prior have exactly sparse inverse kernel matrices, enabling efficient regression and interpolation. The computation of the inverse kernel matrix is reduced from cubic to linear time, and interpolation can be computed in constant time. When using this prior with a linear measurement model, they show exact equivalence to discrete-time smoothing with identical computational complexity when evaluated at the measurement times [35, 24]. Solving the SLAM problem using this method was called Simultaneous Trajectory Estimation And Mapping (STEAM), with Anderson and Barfoot [61] extending the approach to poses using local LTV-SDEs (Section III-C1). Nonlinear models are supported in these works through linearization and iterative optimization.
@@ IV-C2 Optimization
Exploiting the sparsity of the formulation, the STEAM problem can be modeled as a factor graph with binary GP prior factors connecting adjacent states [35, 61], enabling batch [61] and sliding-window optimization [220]. Yan et al. [194, 195] used this to develop a GP version of the incremental optimizer iSAM2 [230], and with iGPMP2 [60] applied this incremental GP approach to motion-planning problems. Sparse interpolated measurement factors allow measurements to be added between the explicitly estimated states [49, 35, 195, 36], and can be readily used within an incremental Bayes tree optimization scheme [195]. Dong et al. [62, 63] provided a formulation for general matrix Lie groups before integrating it into their Simultaneous Trajectory Estimation And Planning (STEAP) framework [197, 198]. Exploring a different direction, Barfoot et al. [205] propose Exactly Sparse Gaussian Variational Inference (ESGVI) to fit a Gaussian to the full Bayesian posterior of the problem states. This approximation ought to be more representative than the Laplace approximation from MAP estimation (Equation 6). They devise a Newton-style iterative solver that avoids the need for analytic derivatives through cubature rules and exploits the inherent sparsity of the problem structure to achieve the same computational complexity as MAP estimation. Wong et al. [206] then extend this approach for hyperparameter learning with an Expectation Maximization (EM) optimization framework.
@@ IV-C3 Applications using the WNOA Prior
The GN TGP approach [35, 61, 62] has been applied in the context of teach and repeat, such as in Visual Odometry (VO) for UAVs [201, 202], or LiDAR and RADAR odometry [216, 224]. Liu et al. [217] and Wang and Gammell [218] have proposed incremental event-based VO pipelines based on this CT representation, employing the GN algorithm in the underlying optimization. Li et al. [212] demonstrate LiDAR-inertial extrinsic calibration and localization in structured environments. Incremental Bayes-tree-based optimizers [230] have been used in SLAM for agricultural monitoring [196] and inertial-GNSS odometry [214]. Recent works [215, 219] have applied TGPs to FMCW LiDAR odometry. Rather than aggregating points (e.g., for scan registration), Wu et al. [215] formulate point-to-plane and radial-velocity residuals for each point. Yoon et al. [219] combine radial-velocity residuals with a gyroscope for linear estimation of the total DoF ego-velocity, which they integrate into poses. Judd and Gammell [204, 203] address the multimotion estimation problem with TGPs in their occlusion-aware VO pipeline, estimating for the ego-motion and trajectories of points on dynamic objects.
@@ IV-C4 Applications using Other Priors
The aforementioned works all use the WNOA prior. Tang et al. [64] propose the ‘constant-acceleration’ White-Noise-On-Jerk (WNOJ) model instead. They achieve higher accuracies in some cases and prove that when the chosen prior does not match the motion, the state estimate will be biased along certain DoF. The WNOJ prior has since been used for online GNSS multi-sensor odometry [213, 227] and target-based multi-sensor extrinsic calibration [208]. Recently, Zheng and Zhu [223] propose multi-sensor LIO using WNOJ for position, WNOA for orientation, and White-Noise-On-Velocity (WNOV) (a.k.a. ‘random-walk’) for IMU biases, and achieve state-of-the-art results. As in Wu et al. [215] and their previous work [30], they avoid measurement aggregation (e.g., LiDAR registration, IMU pre-integration) while retaining real-time performance through the split pose representation and efficient map management. Burnett et al. [225] also utilize the WNOV prior for IMU biases, demonstrating effective LiDAR-inertial and RADAR-inertial odometry on various datasets. Judd and Gammell [210, 211] compare the WNOJ and WNOA priors in their Multimotion Visual Odometry (MVO) pipeline [203, 204]. Johnson et al. [22] recently were the first to compare TGPs and B-splines in visual-inertial localization experiments. They compared WNOJ TGPs to cubic and quintic B-splines, both with and without motion prior factors, and for both the split and joint pose representations. They conclude that with the same motion models and measurements, and when the degree-of-differentiability matches, temporal splines and GPs are similar in accuracy and solve time. Furthermore, they show that motion prior factors act as regularizers, preventing overfitting to noisy measurements. However, these benefits are significantly diminished when an IMU is available. Zheng and Zhu [30] and Johnson et al. [22] investigate the efficacy of the split pose representation and conclude its primary benefit is the reduction of computation time, both at solve and inference time. From the existing literature, Johnson et al. [22] provide insights into computation timings of these CT method variants, while Cioffi et al. [19] demonstrate that CT solve times can be on par with DT. Most recently, Nguyen et al. [65] released their split pose WNOJ implementation and provide localization and calibration examples for a variety of sensor combinations. They further provide a generalization of White-Noise-On-Derivative (WNOD) priors for any order and analytic Jacobians for their residuals. Wong et al. [207] investigate the Singer prior, which represents latent accelerations as a GP with a Matérn kernel. They argue that the data-driven optimization of an additional parameter enables a better representation of the robot’s motion. Burnett et al. [228] utilize the Singer prior for LIO, and comparatively analyze the use of IMU as a motion model input (for preintegration) and as a measurement.